# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### v2

@@ II.2 Definitions of random circuit ensemble and random outcome ensemble
Indeed, this circuit ensemble is similar to the random circuit ensemble employed by the previous random circuit sampling proposals [40, 41, 28, 42, 43] which employ gate-wise Haar random unitary circuits. Next, we introduce our second random circuit ensemble, which is the random permutation circuit ensemble over the fixed shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
@@
We define $\mathcal{P}$ as the uniform distribution over $M$-mode permutation circuits over the fixed Kaleidoscope circuit architecture $\mathcal{B}\mathcal{B}^{*}$ composed of $M\log M$ number of two-mode permutation gates $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$. Each $\bm{P}\sim\mathcal{P}$ represents an $M$ by $M$ permutation matrix, uniformly chosen from $M!$ number of possible permutation matrices.
Note that this random permutation circuit ensemble can be implemented in the shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$ by Lemma 1. Before proceeding, we briefly remark on the motivation for including the permutation circuit ensemble in our random circuit construction. Specifically, for any $M$-mode circuit architecture $\mathcal{A}$, including $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ considered in this work, a circuit drawn from the product circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ guarantees bosonic birthday paradox. The bosonic birthday paradox ensures an upper bound on the probability of obtaining collision outcomes over all possible outcomes of Boson Sampling (see Appendix B for more details). More precisely, for a randomly chosen circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$, the total probability weight of collision outcomes in the Boson Sampling output distribution is upper bounded by $O(\frac{N^{2}}{M})$. Indeed, this bound becomes sufficiently small in the dilute regime $M=\Omega(N^{2})$, which is the primary regime of interest in this work. Accordingly, the bosonic birthday paradox ensures that collision-free outcomes dominate the probability weight for our setup, and thus enables us to confine the outcome space only to collision-free outcomes. Also, it is worth emphasizing that, from an experimental perspective, the setup for our random circuit ensemble can be realized by constructing the local random circuit $\mathcal{H}_{\mathcal{A}}$ and preparing a random input configuration over the experiments.
@@
We define $\,\mathcal{G}_{M,N}$ as the uniform distribution over $\binom{M}{N}$ number of collision-free outcomes of Boson Sampling with $M$ modes and $N$ photons. Each outcome $\bm{s}\sim\mathcal{G}_{M,N}$ is an $M$-dimensional output configuration vector representing the collision-free outcome, such that $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}s_{i}=N$.
III Worst-case hardness of output probability estimation
@@
Given $M=\Omega(N)$, for a fixed collision-free outcome $\bm{s}_{0}$, approximating an output probability $p_{\bm{s}_{0}}(C_{0})$ to within additive imprecision $2^{-O(N)}$ for any $C_{0}$ over linear optical circuit architecture $\mathcal{B}\mathcal{B}^{*}$ is #$\rm{P}$-hard in the worst case.
We briefly sketch the proof of our worst-case hardness result for the shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$. The proof is based on the result by [33], which showed the simulation hardness of exact Boson Sampling with constant-depth linear optical circuits. Specifically, as shown in [33], there exists a constant-depth linear optical circuit that can simulate an arbitrary given quantum circuit using post-selection. To extend this result to our case, we show that the constant-depth worst-case circuit in [33] can be embedded in our circuit architecture $\mathcal{B}\mathcal{B}^{*}$. Combining these arguments, additively approximating the output probability of any quantum circuit can be reduced to additively approximating the output probability of any circuit in $\mathcal{B}\mathcal{B}^{*}$, with imprecision blowup up to the inverse of post-selection probability. Using the fact that the additive approximation of any quantum circuit is #P-hard for certain additive imprecision [42], we can obtain the worst-case hardness of our shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
@@
IV Average-case hardness of output probability estimation
@@
[…]
@@
The following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input an $O(\log N)$-depth random circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$ and a random collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the output probability $p_{\bm{s}}(U)$ within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
@@
We sketch the proof of Theorem 4 in the following by briefly describing the worst-to-average-case reduction process; we leave in Appendix E a detailed proof of Theorem 4. Now, we clarify how to establish the reduction: The goal is to well-estimate worst-case output probability $p_{\bm{s}_{0}}(C_{0})$ for a fixed collision-free outcome $\bm{s}_{0}$ and worst-case circuit $C_{0}$ in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$ (Theorem 3), given oracle access to well-estimated values $p_{\bm{s}}(U)$ for average-case collision-free outcomes over $\bm{s}\sim\mathcal{G}_{M,N}$ and average-case circuits over $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$ (in other words, $U$ is a randomly chosen circuit over $(\mathcal{B}\mathcal{B}^{*})^{q\geq 4}$).
@@
First, since our average-case argument regards both outcomes and circuits, we begin by describing how to fix the outcome in establishing the worst-to-average-case reduction. The remaining problem is then to establish a “circuit-level” worst-to-average-case reduction, from a worst-case circuit to randomly chosen average-case circuits for a fixed output probability, similarly to previous hardness results of random circuit sampling [40, 41, 28, 42, 43]. To fix the outcome, our strategy is to randomly permute both a given worst-case circuit $C_{0}$ in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$ and a given collision-free outcome $s_{0}$. That is, we sample random $M$-mode permutation $\bm{P}_{0}$ uniformly, and permute the worst-case circuit $C_{0}$ and the fixed outcome $\bm{s}_{0}$ in Theorem 3 equally with $\bm{P}_{0}$. Then, the permuted outcome $\bm{s}=\bm{P}_{0}\bm{s}_{0}$ essentially follows the random outcome ensemble $\bm{s}\sim\mathcal{G}_{M,N}$, and the permuted circuit $\bm{P}_{0}C_{0}$ is in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ by Lemma 1. Note that the fixed worst-case output probability $p_{\bm{s}_{0}}(C_{0})$ is now equivalent to $p_{\bm{s}}(\bm{P}_{0}C_{0})$, and thus our new goal is to estimate $p_{\bm{s}}(\bm{P}_{0}C_{0})$ value, given estimated average-case values of $p_{\bm{s}}(U)$.
@@
Next, we describe how to estimate $p_{\bm{s}}(\bm{P}_{0}C_{0})$ given average-case estimation values $p_{\bm{s}}(U)$ for $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$. For the permutation part $\mathcal{P}$ in the average-case circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$, we newly sample an $M$-mode random permutation $\bm{P}_{1}\sim\mathcal{P}$ again, and using the sampled $\bm{P}_{1}$ we set the circuit $\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}$ as a “revised” worst-case circuit in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$. For the local random circuit part $\mathcal{H}_{\mathcal{A}}$ in the average-case circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$, we sample a random circuit $V$ from $\mathcal{H}_{\mathcal{A}}$ and perturb each gate of $V$ by each gate of $\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}$ parameterized by a constant $\theta\in[0,1]$ using Cayley transform method, and obtain a perturbed circuit $V(\theta)$ (see Appendix D for details). Here, by Cayley transform, $V(\theta)$ lies on a continuous path over unitary matrices interpolating $V(0)$ and $V(1)$, where $\theta=0$ corresponds to the random circuit drawn from local random circuit ensemble $V(0)\sim\mathcal{H}_{(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}}$ and $\theta=1$ corresponds to the worst-case circuit $V(1)=\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}$. Notably, $V(0)\bm{P}_{1}$ follows the average-case circuit distribution $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$, and thus, one can expect that the distribution for $V(\theta)\bm{P}_{1}$ for sufficiently small $\theta$ would not be largely deviated from the average-case circuit distribution. Accordingly, one can obtain well-estimated values of $p_{\bm{s}}(V(\theta)\bm{P}_{1})$ for sufficiently small $\theta$ with high probability, given oracle access to the average-case output probabilities (i.e., Theorem 4).
@@
$p_{\bm{s}_{0}}(C_{0})=p_{\bm{s}}(\bm{P}_{0}C_{0})=p_{\bm{s}}(\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}\bm{P}_{1})\quad\xleftarrow[\begin{subarray}{c}V\rightarrow\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}\\ \text{Cayley transform}\\ \text{and}\\ \text{Polynomial interpolation}\end{subarray}]{}\quad p_{\bm{s}}(V\bm{P}_{1})=p_{\bm{s}}(U)\approx\mathcal{O}(\bm{s},U),$ (2)
@@
[…]
So far, we have outlined the proof of Theorem 4, describing how to establish worst-to-average-case reduction. In Appendix D, we introduce in detail the Cayley transform, the circuit perturbation method for the worst-to-average-case reduction in our work. Subsequently, in Appendix E, we provide a detailed proof of Theorem 4, including the imprecision analysis we have missed out in the main text (i.e., why the average-case imprecision level $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$ is necessary to achieve the worst-case imprecision level $2^{-O(N)}$).
@@
See Appendix E
It is worth remarking that, although we have concentrated on the dilute regime such that $\gamma\geq 2$, our average-case hardness result does not rule out the case $21\leq\gamma<$ 2 because it is derived from the worst-to-average-case reduction and our worst-case hardness result allows $M=\Omega(N)$. However, since our average-case hardness result only considers collision-free outcomes, it does not directly lead to the sampling hardness result (i.e., quantum advantage claim for shallow-depth Boson Sampling) by only increasing the allowed additive imprecision for the average-case hardness because collision-free outcomes are expected to occupy a small portion of outcomes when $21\leq\gamma<$ 2. Therefore, to lead to the sampling hardness argument when $21\leq\gamma<$ 2, our average-case hardness result should be extended to also contain collision outcomes for the average-case outcome instances (i.e., instead of $\mathcal{G}_{M,N}$, now uniformly choose over all possible outcomes). For a more complete analysis, in Appendix G, we present a promising approach that can be extended to the average-case hardness result that includes collision outcomes.
V Average-case hardness implies classical simulation hardness
As we have previously discussed, since our average-case hardness result considers both the random outcomes and random circuits, it is not straightforward to show the classical simulation hardness of shallow-depth Boson Sampling as in the original Boson Sampling proposal [1]. Therefore, in this section, we provide a self-contained analysis of how our average-case hardness result leads to the classical simulation hardness arguments of shallow-depth Boson Sampling. Specifically, we show that if the allowed additive error in Theorem 4 for the hardness is improved to a certain imprecision level, an efficient classical algorithm that can approximately simulate the shallow-depth Boson Sampling is unlikely to exist. This emphasizes that improving the imprecision level of the average-case hardness in Theorem 4 is a crucial step for the classical intractability of shallow-depth Boson Sampling.
Similarly to Refs. [1, 40], we define an approximate boson sampler as follows.

### v3

@@ 2.2 Definitions for random circuit and outcome ensemble
Indeed, this circuit ensemble is similar to the random circuit ensemble employed by the previous random circuit sampling proposals [bouland2019complexity, movassagh2023hardness, bouland2022noise, kondo2022quantum, krovi2022average] which employ gate-wise Haar random unitary circuits. Next, we introduce our second random circuit ensemble, which is the random permutation circuit ensemble over the fixed shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
@@
We define $\mathcal{P}$ as the uniform distribution over $M$-mode permutation circuits over the fixed Kaleidoscope circuit architecture $\mathcal{B}\mathcal{B}^{*}$ composed of $M\log M$ number of two-mode permutation gates $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$. Each $\bm{P}\sim\mathcal{P}$ represents an $M$ by $M$ permutation matrix, uniformly chosen from all possible permutation matrices.
Note that this random permutation circuit ensemble can be implemented in the shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$ by Lemma 1. Before proceeding, we briefly remark on the motivation for including the permutation circuit ensemble in our random circuit construction. Specifically, for any $M$-mode circuit architecture $\mathcal{A}$, including $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ considered in this work, a circuit drawn from the product circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ guarantees the bosonic birthday paradox [aaronson2011computational, arkhipov2012bosonic]. The bosonic birthday paradox ensures that the probability of obtaining only collision outcomes is sufficiently large given that the photon number NN is sufficiently smaller than the mode number $M$, as stated in the following lemma.
Lemma 2 (Bosonic birthday paradox for the product circuit ensemble).
Let $M\geq 2N^{2}$, and let $B_{M,N}$ be a set of collision outcomes for $M$ modes and NN photons. Then, for a randomly chosen unitary circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$, the probability to obtain collision outcomes is upper bounded by
The proof of Lemma 2 is provided in Appendix B. Accordingly, the total probability weight of collision outcomes in the Boson Sampling output distribution can be made sufficiently small in the dilute regime $M=\Omega(N^{2})$, which is the primary regime of interest in this work. Therefore, when using the circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$, the bosonic birthday paradox in Lemma 2 ensures that collision-free outcomes dominate the probability weight, thereby enabling us to confine the outcome space only to collision-free outcomes.
It is worth emphasizing that, for the bosonic birthday paradox, the circuit ensemble need not be restricted to $\mathcal{H}_{\mathcal{A}}$. In fact, regardless of the underlying circuit distribution, the initial random permutation layer $\mathcal{P}$ prior to the circuit distribution ensures the bosonic birthday paradox. From an experimental perspective, rather than physically implementing the permutation circuit $\mathcal{P}$, one may instead prepare a uniformly random input configuration, which is operationally equivalent.
Remark 1 (Conditions for the bosonic birthday paradox).
The bosonic birthday paradox is ensured in the dilute regime as long as the input collision-free configuration is prepared uniformly random by any means. This observation opens the possibility of improving hardness results for shallow-depth Boson Sampling by considering more general circuit ensembles beyond $\mathcal{H}_{\mathcal{A}}$.
@@
We define $\mathcal{G}_{M,N}$ as the uniform distribution over $\binom{M}{N}$ number of collision-free outcomes of Boson Sampling with $M$ modes and NN photons. Each outcome $\bm{s}\sim\mathcal{G}_{M,N}$ is an $M$-dimensional output configuration vector representing the collision-free outcome, such that $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}s_{i}=N$.
3 Worst-case hardness of output probability estimation
@@
Given $M=\Omega(N)$, for a fixed collision-free outcome $\bm{s}_{0}$, approximating an output probability $p_{\bm{s}_{0}}(C_{0})$ to within additive imprecision $2^{-O(N)}$ for any $C_{0}$ over linear optical circuit architecture $\mathcal{B}\mathcal{B}^{*}$ is $\rm{P}$-hard in the worst case.
We briefly sketch the proof of our worst-case hardness result for the shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$. The proof is based on the result by [brod2015complexity], which showed the simulation hardness of exact Boson Sampling with constant-depth linear optical circuits. Specifically, as shown in [brod2015complexity], based on measurement-based quantum computation (MBQC) scheme, there exists a constant-depth linear optical circuit that can simulate an arbitrary given quantum circuit using post-selection. To extend this result to our case, we show that the constant-depth worst-case MBQC circuit in [brod2015complexity] can be embedded in our circuit architecture $\mathcal{B}\mathcal{B}^{*}$. Combining these arguments, additively approximating the output probability of any quantum circuit can be reduced to additively approximating the output probability of any circuit in $\mathcal{B}\mathcal{B}^{*}$, with imprecision blowup up to the inverse of post-selection probability. Using the fact that the additive approximation of any quantum circuit is #P-hard for certain additive imprecision [kondo2022quantum], we can obtain the worst-case hardness of our shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
@@
We conclude this section by noting that, it is not necessary to derive worst-case hardness of shallow-depth Boson Sampling exclusively for the circuit architecture $\mathcal{BB}^{*}$. That is, embedding the constant-depth worst-case MBQC circuit [brod2015complexity] into the logarithmic-depth $\mathcal{BB}^{*}$ circuit, as done in our proof of Theorem 3, is not essential for the worst-case hardness of shallow-depth Boson Sampling in principle (and, as will be discussed later, not even for the worst-to-average-case reduction). Although not mandatory, however, since we later make use of $\mathcal{BB}^{*}$ to implement the global permutation during the worst-to-average-case reduction (due to the absence of hiding), we establish worst-case hardness within $\mathcal{BB}^{*}$ so as to simplify the overall circuit architecture to the repeated structure $(\mathcal{BB}^{*})^{q}$ defined in Definition 3.
4 Average-case hardness of output probability estimation
@@
As introduced, our main strategy for the average-case hardness is the worst-to-average-case reduction, using the previous worst-case hardness result in Theorem 3. In other words, we show that if we can well-estimate the output probability $\it{on\,average}$, we can also well-estimate the worst-case output probability in Theorem 3, which means that well-estimating the average-case output probability is also #P-hard. Here, as we have stated previously, our average-case approximation regards both the randomly chosen outcome and the randomly chosen circuit.
@@
The following problem is $\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input an $O(\log N)$-depth random circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$ and a random collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the output probability $p_{\bm{s}}(U)$ within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
@@
We sketch the proof of Theorem 4 in the following by briefly describing the worst-to-average-case reduction process; we leave in Appendix F a detailed proof of Theorem 4. Now, we clarify how to establish the reduction: The goal is to well-estimate worst-case output probability $p_{\bm{s}_{0}}(C_{0})$ for a fixed collision-free outcome $\bm{s}_{0}$ and worst-case circuit $C_{0}$ in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$ (Theorem 3), given oracle access to well-estimated values $p_{\bm{s}}(U)$ for average-case collision-free outcomes over $\bm{s}\sim\mathcal{G}_{M,N}$ and average-case circuits over $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$ (in other words, $U$ is a randomly chosen circuit over $(\mathcal{B}\mathcal{B}^{*})^{q\geq 4}$).
@@
First, since our average-case argument regards both outcomes and circuits, we begin by describing how to fix the outcome in establishing the worst-to-average-case reduction. The remaining problem is then to establish a “circuit-level" worst-to-average-case reduction, from a worst-case circuit to randomly chosen average-case circuits for a fixed output probability, similarly to previous hardness results of random circuit sampling [bouland2019complexity, movassagh2023hardness, bouland2022noise, kondo2022quantum, krovi2022average]. To fix the outcome, our strategy is to randomly permute both a given worst-case circuit $C_{0}$ in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$ and a given collision-free outcome $s_{0}$. That is, we sample random $M$-mode permutation $\bm{P}_{0}$ uniformly, and permute the worst-case circuit $C_{0}$ and the fixed outcome $\bm{s}_{0}$ in Theorem 3 equally with $\bm{P}_{0}$. Then, the permuted outcome $\bm{s}=\bm{P}_{0}\bm{s}_{0}$ essentially follows the random outcome ensemble $\bm{s}\sim\mathcal{G}_{M,N}$, and the permuted circuit $\bm{P}_{0}C_{0}$ is in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ by Lemma 1. Note that the fixed worst-case output probability $p_{\bm{s}_{0}}(C_{0})$ is now equivalent to $p_{\bm{s}}(\bm{P}_{0}C_{0})$, and thus our new goal is to estimate $p_{\bm{s}}(\bm{P}_{0}C_{0})$ value, given estimated average-case values of $p_{\bm{s}}(U)$.
@@
Next, we describe how to estimate $p_{\bm{s}}(\bm{P}_{0}C_{0})$ given average-case estimation values $p_{\bm{s}}(U)$ for $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$. For the permutation part $\mathcal{P}$ in the average-case circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$, we newly sample an $M$-mode random permutation $\bm{P}_{1}\sim\mathcal{P}$ again, and using the sampled $\bm{P}_{1}$ we set the circuit $\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}$ as a “revised" worst-case circuit in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$. For the local random circuit part $\mathcal{H}_{\mathcal{A}}$ in the average-case circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$, we sample a random circuit $V$ from $\mathcal{H}_{\mathcal{A}}$ and perturb each gate of $V$ by each gate of $\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}$ parameterized by a constant $\theta\in[0,1]$ using Cayley transform method, and obtain a perturbed circuit $V(\theta)$ (see Appendix E for details). Here, by Cayley transform, $V(\theta)$ lies on a continuous path over unitary matrices interpolating $V(0)$ and $V(1)$, where $\theta=0$ corresponds to the random circuit drawn from local random circuit ensemble $V(0)\sim\mathcal{H}_{(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}}$ and $\theta=1$ corresponds to the worst-case circuit $V(1)=\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}$. Notably, $V(0)\bm{P}_{1}$ follows the average-case circuit distribution $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$, and thus, one can expect that the distribution for $V(\theta)\bm{P}_{1}$ for sufficiently small $\theta$ would not be largely deviated from the average-case circuit distribution. Accordingly, one can obtain well-estimated values of $p_{\bm{s}}(V(\theta)\bm{P}_{1})$ for sufficiently small $\theta$ with high probability, given oracle access to the average-case output probabilities (i.e., Theorem 4).
@@
$p_{\bm{s}_{0}}(C_{0})=p_{\bm{s}}(\bm{P}_{0}C_{0})=p_{\bm{s}}(\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}\bm{P}_{1})\;\xleftarrow[\begin{subarray}{c}V\rightarrow\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}\\ \text{Cayley transform}\\ \text{\&}\\ \text{polynomial}\\ \text{interpolation}\end{subarray}]{}\;p_{\bm{s}}(V\bm{P}_{1})=p_{\bm{s}}(U)\approx\mathcal{O}(\bm{s},U),$ (3)
@@
So far, we have outlined the proof of Theorem 4, describing how to establish worst-to-average-case reduction. In Appendix E, we introduce in detail the Cayley transform, the circuit perturbation method for the worst-to-average-case reduction in our work. Subsequently, in Appendix F, we provide a detailed proof of Theorem 4, including the imprecision analysis we have missed out in the main text (i.e., why the average-case imprecision level $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$ is necessary to achieve the worst-case imprecision level $2^{-O(N)}$).
@@
See Appendix F
It is worth mentioning that, although we have concentrated on the dilute regime such that $\gamma\geq 2$, our average-case hardness result does not rule out the case $1\leq\gamma<2$ because it is derived from the worst-to-average-case reduction and our worst-case hardness result allows $M=\Omega(N)$. However, since our average-case hardness result only considers collision-free outcomes, it does not directly lead to the sampling hardness result (i.e., quantum advantage claim for shallow-depth Boson Sampling) by only increasing the allowed additive imprecision for the average-case hardness because collision-free outcomes are expected to occupy a small portion of outcomes when $1\leq\gamma<2$. Therefore, to lead to the sampling hardness argument when $1\leq\gamma<2$, our average-case hardness result should be extended to also contain collision outcomes for the average-case outcome instances (i.e., instead of $\mathcal{G}_{M,N}$, now uniformly choose over all possible outcomes). For a more complete analysis, in Appendix H, we present a promising approach that can be extended to the average-case hardness result that includes collision outcomes.
We also remark again that one could further reduce the circuit depth by taking the worst-case circuit $C_{0}$ as the constant-depth MBQC circuit in [brod2015complexity] directly, without encoding into the logarithmic-depth architecture $(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$ as in Theorem 3. In this case, the overall circuit architecture $\mathcal{A}$ for the average-case hardness consists of the constant-depth circuit (for MBQC) sandwiched between the $\mathcal{BB}^{*}$ architectures used for implementing required permutations (i.e., $\bm{P}_{0}$ and $\bm{P}_{1}$), yielding a slightly shallower circuit architecture than $\mathcal{A}=(\mathcal{BB}^{*})^{q\geq 3}$ used in our case. Nevertheless, since we are already in the logarithmic-depth regime (by implementing permutations with $\mathcal{BB}^{*}$), we adopt $\mathcal{A}=(\mathcal{BB}^{*})^{q\geq 3}$ as our overall circuit architecture for the ensemble $\mathcal{H}_{\mathcal{A}}$ in order to provide a unified description of the circuit architecture required for the hardness of shallow-depth Boson Sampling, as well as a clearer layout for experimental implementations.
5 Classical simulation hardness of shallow-depth boson sampling
As we have previously discussed, since our average-case hardness result considers both the random outcomes and random circuits, it is not straightforward to show the classical simulation hardness of shallow-depth Boson Sampling as in the original Boson Sampling proposal [aaronson2011computational]. Therefore, in this section, we provide a self-contained analysis of how our average-case hardness result leads to the classical simulation hardness arguments of shallow-depth Boson Sampling. Specifically, we show that if the allowed additive error in Theorem 4 for the hardness is improved to a certain imprecision level, an efficient classical algorithm that can approximately simulate the shallow-depth Boson Sampling is unlikely to exist. This emphasizes that improving the imprecision level of the average-case hardness in Theorem 4 is a crucial step for the classical intractability of shallow-depth Boson Sampling.
Similarly to Refs. [aaronson2011computational, bouland2019complexity], we define an approximate boson sampler as follows.